# Learning thin deformable object manipulation with a multi-sensory integrated soft hand

### v1

@@ III-E Observation and action
Formally, the action $a_{t}^{in}$ is defined as $(x_{t},z_{t},\theta_{t},\Omega)$, as shown in Fig. 3C. Here, $(x_{t},z_{t},\theta_{t})$ represents the gripper displacement, representing the relative discrepancy between the current and the desired gripper pose after the slip motion. Specifically, $x_{t}$ and $z_{t}$ are the displacement along axis $\alpha$ and $\gamma$, respectively; $\theta_{t}$ denotes the orientation of the gripper about the axis $\beta$. The parameter $x_{t}$ adjusts the distance of the index finger from the object’s edge. Positioning closer to the edge reduces the required manipulation energy, but being too close or beyond the edge can cause multiple layers to flip up. The parameter $z_{t}$ regulates the pressure applied on the object’s surface to ensure adequate friction during manipulation. The parameter $\theta_{t}$ enables the robot to distribute varying amounts of force between the two fingers, and the appropriate distribution will facilitate separation and grasping. Additionally, $\Omega$ is a Boolean vector that represents either the closing or opening of the pneumatic finger, which is set to true at the end of the episode to close the gripper. More details of the observation and action are in Section S5.
III-F Policy architecture
@@
III-G Reward
The episode length is fixed to one in order to learn successful policy within a minimal number of steps. At the end of an episode, we provide a positive reward 1 for successfully grasping a single layer of an object. In contrast, results where zero, two or more layers are grasped simultaneously receive a reward of 0. The reward $r_{t}$ in both the book and fabric environments is determined automatically with information naturally possessed by the manipulated object. Specifically, rewards in the book environment are determined by referencing the page number and in the fabric environment by matching QR code data to a pre-set sequence in the manufacturer’s database.
III-H Multisensory encoder
@@
$l_{t}^{\kappa}=g_{\kappa}(l_{t}^{\text{vis}},l^{ind}_{t},l^{thu}_{t},l_{t}^{\text{c}},l_{t}^{\text{prop}},l_{t}^{\text{aux}}).$
@@
III-I Training in real-world via self-supervision
We train the manipulation policy in the real world (Fig. 3A). Operating in parallel environments, robotic arms execute these tasks while continuously updating the policies according to rewards. Specifically, we train our model through trial and error with the following procedure:
At the start of each training episode, two robotic arms use their wrist cameras to capture RGB-D images of environments, which are subsequently processed by the slip module to predict the position and direction of the slip motion. Following this prediction, the robot starts to perform a slip motion from its initial posture in each environment to obtain multisensory observations. Then, the robot downloads the latest parameters for both the inner and outer loop policies from the optimizer. The outer loop policy $\pi^{out}$ selects an action space based on multisensory observations. The inner loop policy $\pi^{in}$ explores within the range of the action space to output a concrete action. Rewards are obtained after robot execution. Finally, training episodes generated from both environments are added to the replay buffer, from which the optimizer samples to update the policy. Both the inner and outer loop policies are periodically updated using the Adam optimizer [59]. The robots continuously collect episodes in these environments. Upon reaching the final book page or fabric layer, the system activates a recycling mechanism to reset the environment. Hyperparameters about the policy training are in Tab. S2.
Figure 4: Robust singulation and grasping in challenging scenarios. The policy learned by the presented method successfully overcame a full spectrum of challenging thin, deforamble objects, which include various textures of paper (A to C, E to G), suit fabrics from different seasons (I and J), hotel towels (D), clothing (H), and complex hybrid objects (K and L). Except for the book made of printer paper (A) and winter suit fabric (I), these objects had never been encountered during RL policy training.
@@ III-I Training in real-world via self-supervision
Figure 6: Results of manipulating thin, deformable objects under various ablation methods and experimental conditions. (A to C) Illustration of experiments conducted on workspaces with varying inclinations. (D to E) Performance heatmaps show task success rates for the proposed approach compared to ablation methods, evaluated over three workspace tilt angles. (F) Effects of workspace tilt angles on manipulation performance with ablation methods, presented via a boxplot that pictures the mean success rate achieved by our method and the comparative ablation methods across all experiments. (G) Training curves for manipulation policy learning across different ablation methods.
@@ IV Results
Movie 1 summarizes the proposed approach and the experiments. We conducted policy training in two parallel environments: a book made of printer paper and a booklet crafted from winter suit fabric. These materials, with their vast differing compressive strengths and thicknesses, cover characteristics of a broad range of thin and deformable objects. Such training objects allow robots to experience and learn from the diverse physical properties and behaviors of thin, deformable objects, ensuring the generalization of the learned policy. For all the following experiments, we utilized the same RL policy on the robot without further tuning.
@@ IV-A Robust singulation and grasping across thin deformable objects
We tested our system across a variety of scenarios involving the singulation and grasping of thin, deformable objects. These scenarios include turning book pages (Fig. 4, A to C), emptying paper boxes (Fig. 4, E to G), displaying fabrics (Fig. 4, I and J), and picking hotel towels and clothes (Fig. 4, D and H). Two hundred trials were conducted for each scenario configuration. The success of each attempt is determined by whether the robot picks only a single layer/object. The robot demonstrated consistent robustness across all scenarios (Fig. 6D).
@@ IV-B Gravity-induced challenges
In this section, we explore how our robot singulates and grasps thin, deformable objects under the challenging conditions of different workspace tilt angles (Fig. 5). Our experiments tested three tilt angles: $0\degree$, $30\degree$, and $60\degree$ (Fig. 6, A to C). We manually adjusted the robot’s initial pose at different tilt angles to ensure the workspace was in view of the camera. The measured degree of workspace inclinations was only used to calculate the coordinate transformation used to perform slip motion without affecting the decision of the manipulation policy.
As the tilt angle increases, friction between the object layers diminishes, and layers become fluffier. Our robot could handle these challenging conditions gracefully, retaining robust manipulation (Fig. 6D). The slip module enables the robot to capture changes in the object’s surface and interlayer properties caused by the increased tilt angle with sensors. The learned policy adjusts actions based on multi-sensory inputs, allowing dynamic response to changes in the environment. The robot’s compliance further helps fingers conform to the object and evenly distribute contact force. Qualitative observation of the robot’s behavior shows that the learned policy often tends to apply increased pressure compared to manipulation on a flat surface, thereby enhancing friction between the fingers and the object to maintain stable contact.
The success rate distribution across different tilt angles reveals a decline as the tilt angle increases (Fig. 6G). This decline is attributable to the fact that higher tilt angles push the dynamics and interactions of the thin, deformable objects outside the distribution encountered during training on a flat surface. One way to enhance performance is by including various gravity conditions during training. See Section S9 for experiments on the policy trained simultaneously across various workspace tilt angles. These experiments bring our robot closer to a wide range of situations encountered in real-world scenarios. See Movie 1 about how our robot helps a violinist turn sheet music placed on the stand at an angle.
@@ IV-C Analysis of the underlying mechanisms of the task and the robot
To investigate the mechanics of grasping thin deformable objects across different materials, we recorded a successful grasp motion and repeated the action across various areas of the objects. Specifically, using the upper-right corner of the object as the reference point, we incrementally adjust the $\alpha$-axis and $\beta$-axis offsets in action, which cover a 6x6 cm area. This region was divided into 36 individual 1x1 cm cells, with 100 grasp attempts conducted in each cell. Fig. 7 presents the performance heatmaps for both paper and fabric materials.
For the paper material, the performance heatmap exhibits a distinct pattern with higher success rates concentrated along the diagonal (angle bisector) of the grid. This aligns with the intuitive understanding that turning a book page diagonally requires the least energy. Meanwhile, areas near the edge of the paper show reduced success rates due to the tendency to grasp multiple layers simultaneously. Areas closer to the inner region yield very low success rates (evident in the lower-left corner of the heatmap). This behavior reflects the inherent properties of paper, which has high compressive strength and resistance to bending. When the robot applies force away from the edge, the paper resists deformation, requiring significantly greater force to achieve sufficient bending to turn the page.
For fabric material, the performance heatmap followed a slightly different pattern. While higher success rates were also concentrated along the diagonal, the area of optimal performance shifted farther from the edges compared to paper. Furthermore, fabric material demonstrated relatively consistent performance even as the grasp point moved closer to the inner regions. This behavior can be attributed to the lack of compressive strength for the fabric material. Unlike paper, fabric deforms easily and folds or bends in the direction of the applied force.
These findings reveal the distinct mechanics underlying the grasping of different materials and suggest the necessity of obtaining comprehensive environmental information through multisensory integration. It also further justifies the design of action parameters.
@@ IV-D Evaluating the contribution of soft hand
First, we tested the position-based policy trained on the suit fabric (Tab. I, first row). During testing, each layer of fabric was often compressed due to repeated grasping, resulting in inaccurate initial gripper position and large gaps between the fingers and the object‘s surface. Despite lacking sensors to perceive this, the robot achieved a 47% success rate. In contrast, using a rigid gripper would result in losing contact with the fabric under such conditions. Our soft hand’s passive compliance allows it to accommodate position inaccuracies to a certain extent and maintain continuous contact, which is critical for successful manipulation.
We then evaluated the performance of the position-based policy trained on printed books (Tab. I, second row). The policy demonstrated remarkable effectiveness, achieving an 80% success rate when tested on printed books, a performance close to that of our complete system. The book pages are of uniform thickness, exhibit similar inter-layer interactions, and are difficult to compress. Using a rigid gripper to learn the policy led to collisions and emergency stops during exploration. Similarly, hard-coded actions resulted in collisions and stops due to the robot’s repeatability being less than the page thickness. In contrast, the passive compliance of our soft hand absorbs minor imprecisions in operation, ensuring gentle interactions, preventing page tearing, and achieving successful singulation and grasping.
In addition, we tested the performance of a position-based policy trained using both printed books and suit fabric (Tab. I, third row). The distinct properties of these materials led to inconsistent results when identical actions were applied at the same height. As a result, the robot struggled to learn from the reward signal reliably. Nevertheless, we achieved a certain level of success on both training and novel objects, underscoring the value of passive compliance in adapting to diverse material properties without the need for precise control.
These results show the critical role of passive compliance in successfully singulating and grasping thin, deformable objects. Our soft hand enabled the robot to perform tasks more successfully and robustly, a feat unachievable with rigid grippers. This demonstrates the principle that passive compliance can bridge the gap between precise control and the unpredictable nature of real-world interactions.
@@ IV-E Ablation studies
To evaluate the influence of tactile and F/T sensors, we compared our complete approach, “Ours,” with the ablation methods: “OnlyVision” (OV), which excludes both tactile and F/T sensors during policy learning, and “NoTouch” (NT), which excludes only touch information during policy learning.
The OV method, relying solely on depth observation, exhibits a feature bias that manifests as a heavy preference for height information. The learned policy fails to internalize the principles and mechanisms for singulating and grasping thin, deformable objects, affecting the overall performance and generalization. This limitation was reflected in the low success rate of the OV method across various objects (Fig. 6D). In contrast, both NT and Ours incorporated multimodal sensors, providing a more comprehensive understanding of the environment. This allowed the learned policy to focus on critical mechanisms for task success. Our method, which integrates additional tactile sensors compared to NT, is able to obtain local contact features, further enhancing the robustness and generalization ability of learned policy, as evidenced by its superior performance (Fig. 6D). These findings highlight the importance of multi-modal perception in delivering comprehensive task information, thereby preventing feature dependence and facilitating effective policy learning for singulating and grasping thin, deformable objects.

### v2

@@ III-B2 Action
Formally, the action $a_{t}^{in}$ is defined as $(x_{t},z_{t},\theta_{t},\Omega)$, as shown in Fig. 3C. Here, $(x_{t},z_{t},\theta_{t})$ represents the gripper displacement, representing the relative discrepancy between the current and the desired gripper pose after the slip motion. Specifically, $x_{t}$ and $z_{t}$ are the displacement along axis $\alpha$ and $\gamma$, respectively; $\theta_{t}$ denotes the orientation of the gripper about the axis $\beta$. The parameter $x_{t}$ adjusts the distance of the index finger from the object’s edge. Positioning closer to the edge reduces the required manipulation energy, but being too close or beyond the edge can cause multiple layers to flip up. The parameter $z_{t}$ regulates the pressure applied to the object’s surface to ensure adequate friction during manipulation. The parameter $\theta_{t}$ enables the robot to distribute varying amounts of force between the two fingers, and the appropriate distribution will facilitate separation and grasping. Additionally, $\Omega$ is a Boolean vector that represents either the closing or opening of the pneumatic finger, which is set to true at the end of the episode to close the gripper. More details of the observation and action are in Section S5.
III-B3 Reward
At the end of an episode, we provide a positive reward 1 for successfully grasping a single layer of an object. In contrast, results where zero, two or more layers are grasped simultaneously receive a reward of 0. The reward $r_{t}$ in both the book and fabric environments is determined automatically with information naturally possessed by the manipulated object. Specifically, rewards in the book environment are determined by referencing the page number, and in the fabric environment by matching QR code data to a pre-set sequence in the manufacturer’s database. The episode length is fixed to one in order to learn a successful policy within a minimal number of steps.
III-C Observations via slip module
The slip module serves two primary functions: obtaining task-related sensory information and providing an initial robot pose to minimize unnecessary exploration. To obtain rich multisensory observations, we employ a slip motion to proactively interact with the object (Fig. 2B). This process begins with the robot capturing aligned RGB and depth images from above the object. The RGB image is processed through the Segment Anything Model (SAM) [56] to generate object masks. These binary masks are rotated at various angles and subsequently input into the slip network. The slip network outputs a set of affordance maps, one for each input rotation angle. Each map contains values corresponding to the slip motion primitive, which are parameterized by the pixel’s location and its rotation angle within the map. The pixel with the highest affordance value is selected; its spatial position identifies the slip point, and the rotation angle of the corresponding input mask determines the slip direction.
Once the slip point and direction are determined, the depth image is utilized to ascertain the 3D coordinates of this selected pixel in the camera’s frame. This information, combined with a pre-established hand-eye calibration relating the camera and robot fingertip coordinate systems, allows the robot to move to the slip point.
Next, the robot executes the slip motion. First, the robot’s hand rotates such that the $\alpha$ axis (the line connecting the two fingertips) aligns with the slip direction. The plane composed of the $\alpha$ and $\beta$ axes is parallel to the workspace. We define the workspace as the planar supporting surface upon which objects are placed for manipulation experiments, physically corresponding to the table or platform in the robot’s operational environment. During training, the workspace inclination remains consistently at 0 degrees.
The index finger then moves to the predicted slip point, stopping at a standoff distance of 3 mm above the object’s surface, with this height being estimated using the depth camera feed. Contact is then initiated by applying positive air pressure, causing the soft finger to inflate and bend downwards to slip onto the object’s surface. After 1 second, we record sensory information. The air pressure is then released, and the fingers relax. This 3 mm standoff distance was determined empirically; it provides a balance, being large enough to prevent unintended collisions due to minor sensing inaccuracies or positioning errors, yet small enough to ensure reliable surface contact and the acquisition of high-quality sensory data upon slipping.
After a slip, we record 3D arrays of deformation flows from the tactile sensors as touch observation ($o_{t}^{ind},o_{t}^{thu}$), along with readings from the F/T sensor as the muscle proprioception $o_{t}^{pro}$, including forces $(f{\alpha},f{\beta},f{\gamma})$ and torques $(m{\alpha},m{\beta},m{\gamma})$. The three main axes of the F/T sensor are aligned with the three corresponding axes of the $H_{\alpha\beta\gamma}$. This alignment makes the force and torque data measured by the sensor more intuitive. Moreover, we crop the area surrounding the slip point on the depth image to serve as visual observation $o_{t}^{vis}$.
Through the robot’s purposive movements prior to manipulation, observations are made that maximize the information uptake on the relevant properties of the target object. These observations provide insights into the elaborate interplay between static and dynamic sensory cues, such as texture, softness, and friction. The slip network is pre-trained using synthetic data before the robot learns in the real world. Humans annotated slip points on a few images featuring a book, a shirt, and a spring pancake. We employed data augmentation techniques, including rotation, translation, and scaling, on the object masks of these images to simulate different object sizes and positions. More details of the slip network architecture are in Section S6.
III-D Policy architecture
@@
III-E Multisensory encoder
@@
$l_{t}^{\kappa}=g_{\kappa}(l_{t}^{\text{vis}},l^{ind}_{t},l^{thu}_{t},l_{t}^{\text{c}},l_{t}^{\text{pro}},l_{t}^{\text{aux}}).$ (4)
@@
III-F Training details
The training process of our method consists of two distinct phases: 1) training the slip module using synthetic data, and 2) learning the manipulation policy through reinforcement learning in the real world.
III-F1 Slip module training
The slip module is independently trained through supervised learning prior to the RL policy training phase. We collected images of three representative objects: a printed book, a shirt, and a stack of spring pancakes. For each image, appropriate slip points and corresponding directions were manually annotated. To ensure robust generalization to various object positions and orientations, we performed extensive data augmentation, including random rotations, translations, and scaling transformations, generating a diverse synthetic dataset. The slip network is trained on this augmented dataset using the Adam optimizer with cross-entropy loss. Once trained, the slip network parameters are frozen and subsequently utilized for predicting slip points during the RL policy learning phase.
III-F2 Manipulation policy training via RL
In this phase, robots learn the manipulation policy in the real world (Fig. 3A). Operating in parallel environments, robotic arms execute these tasks while continuously updating the policies according to rewards. Specifically, we train our model through trial and error with the following procedure:
At the start of each training episode, two robotic arms use their wrist cameras to capture RGB-D images of environments, which are subsequently processed by the pretrained (and now frozen) slip module to predict the position and direction of the slip motion. Following this prediction, the robot starts to perform a slip motion from its initial posture in each environment to obtain multisensory observations. Then, the robot downloads the latest parameters for both the inner and outer loop policies from the optimizer. The outer loop policy $\pi^{out}$ selects an action space based on multisensory observations. The inner loop policy $\pi^{in}$ explores within the range of the action space to output a concrete action. Rewards are obtained after robot execution. Finally, training episodes generated from both environments are added to the replay buffer, from which the optimizer samples to update the policy. Both the inner and outer loop policies are periodically updated using the Adam optimizer [59]. The robots continuously collect episodes in these environments. Upon reaching the final book page or fabric layer, the system activates a recycling mechanism to reset the environment. The policy training was completed in approximately 1700 trials across the two training environments. Hyperparameters about the policy training are in Table S3.
Figure 4: Robust singulation and grasping in challenging scenarios. The policy learned by the presented method successfully overcame a full spectrum of challenging thin, deformable objects, which include various textures of paper (A to C, E to G), suit fabrics from different seasons (I and J), hotel towels (D), clothing (H), and complex hybrid objects (K and L). Except for the book made of printer paper (A) and winter suit fabric (I), these objects had never been encountered during RL policy training.
@@ III-F2 Manipulation policy training via RL
Figure 6: Results of manipulating thin, deformable objects. (A to C) Experiments conducted on workspaces with varying inclinations. (D to E) Performance heatmaps display task success rates for the proposed approach in comparison to ablation methods. (F) Effects of workspace tilt angles on manipulation performance with ablation methods, presented via a boxplot. (G) Training curves for manipulation policy learning across different ablation methods.
@@ IV Results
Movie 1 summarizes the proposed approach and the experiments. We conducted policy training in two parallel environments: a book made of printer paper and a booklet crafted from winter suit fabric. These materials, with their vastly differing compressive strengths and thicknesses, cover characteristics of a broad range of thin and deformable objects. Such training objects allow robots to experience and learn from the diverse physical properties and behaviors of thin, deformable objects, ensuring the generalization of the learned policy. For all the following experiments, we utilized the same RL policy on the robot without further tuning (Fig. 4 and Fig. 5).
@@ IV-A Robust singulation and grasping across thin deformable objects
We tested our system across a variety of scenarios involving the singulation and grasping of thin, deformable objects. The thickness of these objects is shown in Table S6. These scenarios include turning book pages (Fig. 4, A to C), emptying paper boxes (Fig. 4, E to G), displaying fabrics (Fig. 4, I and J), and picking hotel towels and clothes (Fig. 4, D and H). Two hundred trials were conducted for each scenario configuration. The success of each attempt is determined by whether the robot picks only a single layer/object. The average duration per grasp is approximately 16 seconds. This includes the entire sequence: moving from the initial pose, executing the slip motion, successfully grasping and turning a single page, and returning to the starting position. The robot demonstrated consistent robustness across all scenarios (Fig. 6D).
@@ IV-B Gravity-induced challenges
In this section, we explore how our robot singulates and grasps thin, deformable objects under the challenging conditions of different workspace tilt angles (Fig. 5). Our experiments tested three tilt angles: $0\degree$, $30\degree$, and $60\degree$ (Fig. 6, A to C). We manually adjusted the robot’s initial pose at different tilt angles to ensure the workspace was in view of the camera. The measured degree of workspace inclinations was only used to calculate the coordinate transformation used to perform slip motion, without affecting the decision of the manipulation policy.
As the tilt angle increases, friction between the object layers diminishes, and layers become fluffier. Our robot could handle these challenging conditions gracefully, retaining robust manipulation (Fig. 6D). The slip module enables the robot to capture changes in the object’s surface and interlayer properties caused by the increased tilt angle with sensors. The learned policy adjusts actions based on multi-sensory inputs, allowing dynamic responses to changes in the environment. The robot’s compliance further helps fingers conform to the object and evenly distribute contact force. Qualitative observation of the robot’s behavior shows that the learned policy often tends to apply increased pressure compared to manipulation on a flat surface, thereby enhancing friction between the fingers and the object to maintain stable contact.
The success rate distribution across different tilt angles reveals a decline as the tilt angle increases (Fig. 6G). This decline is attributable to the fact that higher tilt angles push the dynamics and interactions of the thin, deformable objects outside the distribution encountered during training on a flat surface. These experiments bring our robot closer to a wide range of situations encountered in real-world scenarios. See Movie 1 about how our robot helps a violinist turn sheet music placed on the stand at an angle.
@@ IV-C Analysis of the underlying mechanisms of the task and the robot
To investigate the mechanics of grasping thin deformable objects across different materials, we recorded a successful grasp action, and this fixed action was applied across different regions of the object. Specifically, using the upper-right corner of the object as the reference point, we incrementally adjust the $\alpha$-axis and $\beta$-axis offsets, which cover a 6x6 cm area. This region was divided into 36 individual 1x1 cm cells, and the recorded grasp action was executed 100 times at each cell. Fig. 7 presents the performance heatmaps for both paper and fabric materials.
For the paper material, the performance heatmap exhibits a distinct pattern with higher success rates concentrated along the diagonal (angle bisector) of the grid. This aligns with the intuitive understanding that turning a book page diagonally requires the least energy. Meanwhile, areas closer to the inner region yield very low success rates (evident in the lower-left corner of the heatmap). This behavior reflects the inherent properties of paper, which has high compressive strength and resistance to bending. When the robot applies force away from the edge, the paper resists deformation, requiring significantly greater force to achieve sufficient bending to turn the page.
For fabric, the pattern differs. While higher success rates were also concentrated along the diagonal, the area of optimal performance shifted further from the edge. And the performance remains more consistent when moved closer to the inner regions. This behavior can be attributed to the lack of compressive strength of the fabric material, making it easier to fold or bend in the direction of the applied force.
Most importantly, this experiment reveals the critical role of our learned manipulation policy. The different distributions of grasping success rates on different materials indicate that the success cannot be determined solely by the initial pose. Instead, it depends heavily on the robot’s ability to adaptively select the appropriate grasping action based on the sensory input after the slip module provides the initial pose. In this process, the slip module collects rich multi-sensory information. This sensory information captures detailed environmental states, such as material properties, interlayer friction. Our reinforcement learning policy then uses this multisensory input to infer and execute the optimal action for successful singulation and grasping in each specific scenario.
@@ IV-D Evaluating the contribution of soft hand
First, we tested the position-based policy trained on the suit fabric (Table I, first row). During testing, each layer of fabric was often compressed due to repeated grasping, resulting in inaccurate initial gripper position and large gaps between the fingers and the object‘s surface. Despite lacking sensors to perceive this, the robot achieved a 47% success rate. In contrast, using a rigid gripper would result in losing contact with the fabric under such conditions. Our soft hand’s passive compliance allows it to accommodate position inaccuracies to a certain extent and maintain continuous contact, which is critical for successful manipulation.
We then evaluated the performance of the position-based policy trained on printed books (Table I, second row). The policy demonstrated remarkable effectiveness, achieving an 80% success rate when tested on printed books, a performance close to that of our complete system. The book pages are of uniform thickness, exhibit similar inter-layer interactions, and are difficult to compress. Using a rigid gripper to learn the policy led to collisions and emergency stops during exploration. Similarly, hard-coded actions resulted in collisions and stops due to the robot’s repeatability being less than the page thickness. In contrast, the passive compliance of our soft hand absorbs minor imprecisions in operation, ensuring gentle interactions, preventing page tearing, and achieving successful singulation and grasping.
In addition, we tested the performance of a position-based policy trained using both printed books and suit fabric (Table I, third row). The distinct properties of these materials led to inconsistent results when identical actions were applied at the same height. As a result, the robot struggled to learn from the reward signal reliably. Nevertheless, we achieved a certain level of success on both training and novel objects, underscoring the value of passive compliance in adapting to diverse material properties without the need for precise control.
These results show the critical role of passive compliance in successfully singulating and grasping thin, deformable objects. Although impedance control with rigid robotic hands can also be a viable method, it faces several challenging limitations. These include a strong reliance on accurate modeling of contact dynamics, increased control complexity due to the need to specify additional parameters (e.g., desired force trajectories), and significant safety concerns during real-world exploration. Our soft hand enabled the robot to perform tasks more successfully and robustly, a feat unachievable with rigid grippers. This demonstrates the principle that passive compliance can bridge the gap between precise control and the unpredictable nature of real-world interactions.
@@ IV-E Ablation studies
To evaluate the influence of tactile and F/T sensors, we compared our complete approach, “Ours,” with the ablation methods: “OnlyVision” (OV), which excludes both tactile and F/T sensors during policy learning, “NoForce” (NF), which excludes F/T sensors during policy learning, and “NoTouch” (NT), which excludes only touch information during policy learning.
The OV method, relying solely on depth observation, exhibits a feature bias that manifests as a heavy preference for height information. The learned policy fails to internalize the principles and mechanisms for singulating and grasping thin, deformable objects, affecting the overall performance and generalization. This limitation was reflected in the low success rate of the OV method across various objects (Fig. 6D).In contrast, both NF, NT, and Ours incorporated multimodal sensors, providing a more comprehensive understanding of the environment. This allowed the learned policy to focus on critical mechanisms for task success. Quantitatively, both NT and NF underperform relative to Our method. While the tactile sensors in the NF are effective at capturing local contact events, they provide only a localized view of the contact surface. In contrast, the F/T sensor in the NT configuration provides a global proprioceptive signal that reflects distributed contact forces and torques across the entire hand-object interface, particularly those not captured directly by the fingertips. These global cues could contain critical information for inferring macroscopic object states but lack local material information. Our method integrates both modalities, thus leveraging the complementary strengths of each, as evidenced by its superior performance (Fig. 6D). These findings highlight the importance of multi-modal perception in delivering comprehensive task information, thereby preventing feature dependence and facilitating effective policy learning for singulating and grasping thin, deformable objects.